OpenStudy (anonymous):

cos(x+(pi/2))=-sinx verify the identity

OpenStudy (anonymous):

\[\cos(x+(\pi/2))=-sinx\]

OpenStudy (anonymous):

Recall that \[\cos(A+B) = cosAcosB - sinAsinB\]

OpenStudy (anonymous):

so solve for \[= cosxcos \frac{ \pi }{ 2 } - sinxsin \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

so -sinx

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep that's right

OpenStudy (anonymous):

alright great thank u !! can you help me with one more?

OpenStudy (anonymous):

post it... i'll be around

OpenStudy (anonymous):

alright